# Viscous-inertial waves on the surface of the Sun: modeling, forward and inverse problems

### v2

@@ 5.1 Full measurement
With TCC, we further obtain unique identifiability of $\Omega$ given $\gamma$ and vice versa. To this end, we denote by $\mathrm{Arg}(z)\in[0,2\pi)$ the principal value of the argument of a complex number $z$.
@@
    $\Delta_{m}\Psi\neq 0\text{ a.e.~on }I,\quad\mathrm{Arg}(\Psi)-\mathrm{Arg}(\Delta_{m}\Psi)\notin\{-\pi,0,\pi\}\text{ a.e.~on }I.$ (44)
@@
From this, along with the characterization of the solution set given the TCC in [16, Proposition 2.1], we deduce $\Omega^{*}-\Omega^{\dagger}\in\mathcal{N}(F^{\prime}(\Omega^{\dagger}))=\{0\}$ for any solution $\Omega^{*}$, $\Omega^{\dagger}$ of the inverse problem. We thus claim the unique identifiability of $\Omega$ given full measurement.
@@ 5.2 Partial measurement including Cauchy data
We now turn to the case of restricted measurements, where observations are limited to $I^{\prime}=(\epsilon,\pi-\epsilon)$ as described in Section 2.3. As a simplification, we additionally assume that Cauchy data at the boundary of $I^{\prime}$ can be measured exactly.
We follow the analysis for the full Sun, while specifically leveraging the well-posedness and regularity results for the restricted domain established in Section 3.5, Proposition 9.
@@
$\Psi|_{\partial I^{\prime}}=\Psi^{\dagger}|_{\partial I^{\prime}}\quad\text{and}\quad\Psi^{\prime}|_{\partial I^{\prime}}=(\Psi^{\dagger})^{\prime}|_{\partial I^{\prime}},$ (46)
@@
It is worth remarking that the constraint assumption in (46), which requires knowledge of the exact solution on the boundary, is an essential component to the proof of Theorem 17; relaxing this condition poses an interesting question for future research.
@@ 6 Numerical experiments
We implemented the accelerated Nesterov-Landweber method to simultaneously reconstruct the scalar viscosity and latitudinal differential rotation parameters, $(\gamma,\Omega)$. A convergence analysis of this method has been carried out in [19, 29] under the tangential cone condition, which we have been verified for our problem in Section 5. The Nesterov-Landweber iteration proceeds as follows: initiate $p_{-1}=p_{0}:=(\gamma_{\mathrm{int}},\Omega_{\mathrm{int}})$, then run
$z_{k}$ $=p_{k}+\frac{k-1}{k+\alpha-1}\left(p_{k}-p_{k-1}\right)\qquad$ $\alpha=3$ (47)
@@
Here, the forward map $F$ corresponds to the wave equation at single frequency $\omega$ and longitudinal wavenumber $m$. We allow for different observation strategies: full data, leaked data, and real part measurements. The Hilbert space adjoints $F^{\prime}(z_{k})^{*}$ are derived in Section 4.2. The inversion takes place in the following function spaces: $\gamma\in\mathbb{R}$, $\Omega\in H^{2}({r{\mathbb{S}^{2}}},m=0)$, thus ensuring convergence of the iterative reconstruction.
The numerical wave solution uses a fourth-order finite difference scheme [39] on a uniform grid of 100 points spanning the latitude range $(0,\pi)$. To generate ground truths while avoiding inverse crimes, we explicitly choose analytic $\Psi_{m}$$\gamma$ and $\Omega$ with Neumann (Fig. 1) or Dirichlet (Fig. 2) boundary, then explicitly obtain the corresponding source $f$. The discrete measurement $\underline{y}^{\delta}$ is corrupted by independent additive complex Gaussian noise with variance $\sigma^{2}$ on 100 equidistant measurement points and $\sigma$ chosen to have relative noise levels $\|\underline{y}^{\delta}-\underline{y}\|_{2}/\|\underline{y}\|_{2}\in\{0.01,0.05,0.2\}$.
We initialize the algorithm with $\gamma_{\mathrm{int}}=-3\gamma_{\mathrm{true}}$ and $\Omega_{\mathrm{int}}=0$, corresponding to no prior knowledge of the ground truth, and iterate with a step size $\mu_{k}$ determined by backtracking line search. The stopping criterion is the first instance when the residual falls below the threshold $\tau\delta$ (with $\tau=1.03$), which in practice yields better results than choosing $\tau>2$ [19]. The algorithm typically completes 200 iterations within one second on an i7-1255U CPU (4.70 GHz).
@@
Figure 1, row 1 shows the simultaneous recovery of $(\gamma,\Omega)$ from full measurement of the state $\Psi$ at $(\omega,m)=(3,3)$ and with $1\%$ data noise. Despite of an initial guess far from the ground truth, the reconstruction closely approximates the exact parameters. The reconstructed wave field also matches the true state in both real and imaginary parts.
@@
Figure 1, rows 2–3 present the practical case where no data are available near the poles. In our numerical experiments, we did not impose additional Cauchy boundary data in Theorem 17, and simply computed the adjoint as described in Remark 13. With 20% of the data leaked, the reconstruction quality remains nearly as high as with full data. For 50% data loss, the recovery of $\Omega$ moderately deteriorates. Nonetheless, the algorithm exhibits robust performance, consistent with our convergence results in Section 5.2.
@@
Figure 1, row 4 considers the scenario where only the real part of $\Psi$ is measured. Compared to row 3 (which also has 50% data leakage, but full complex data), the absence of the imaginary component further impairs the reconstruction, most notably in missing the secondary peak of the differential rotation.
@@
Finally, Figure 2 reports the results for $(\omega,m)=(1,2)$ with 20 % data leakage and different measurement noise levels. At 1% noise, the recovered parameters remain very close to the truth, as evidenced by small relative errors ($\||p-p_{\text{true}}\|_{L^{2}}/\|p_{\text{true}}\|_{L^{2}}$). At 5% noise, reconstruction quality is slightly reduced, while at 20% noise, the outcome degrades significantly.
@@ 7 Conclusions and Outlook
In this paper, we examined the modeling, as well as the associated forward and inverse problems, for inertial waves on the surface of the Sun. Formulating the dynamics in terms of a stream function leads to a fourth-order elliptic differential equation on the sphere. We rigorously proved well-posedness of this equation, including explicit smallness conditions for uniqueness, as well as conditions for non-uniqueness via analytic Fredholm theory, both in agreement with empirical findings in helioseismology. Our studies represent the starting point for further theoretical and numerical studies of resonances and inertial modes, as well as the associated inverse problems.
@@
Next steps include the incorporation of physical effects such as radial motion, pressure, and gravity [14] into the model and inversions with solar-like parameters, including realistic convection noise [32], as well as real observational data from the Helioseismic and Magnetic Imager. Another promising avenue involves leveraging data-driven model discovery techniques [1, 30] to refine existing theoretical models by learning hidden laws.
@@ Appendix A Differential geometry and function spaces on spheres
independent of the choices of $\tilde{f}$ and $\tilde{\textbf{{f}}}$ (see [28, p. 72–73]). Here $\operatorname{div_{\mathrm{h}}}\textbf{{f}}$ and $\operatorname{curl_{\mathrm{h}}}\textbf{{f}}$ are scalar functions, while $\operatorname{\mathbf{grad}_{\mathrm{h}}}f$ and $\operatorname{\mathbf{curl}_{\mathrm{h}}}v$ are tangential vector fields on $r{\mathbb{S}^{2}}$}. In spherical coordinates, writing $\textbf{{f}}=f_{\theta}\mathbf{e}_{\theta}+f_{\phi}\mathbf{e}_{\phi}$ with $f_{\theta}:=\textbf{{f}}\cdot\mathbf{e}_{\theta}$ and $f_{\phi}:=\textbf{{f}}\cdot\mathbf{e}_{\phi}$, we have
@@
These differential operators satisfy the identities ([28, Thm. 2.5.19]):
$\begin{split}&\operatorname{curl_{\mathrm{h}}}\operatorname{\mathbf{grad}_{\mathrm{h}}}=0,\qquad\operatorname{div_{\mathrm{h}}}\operatorname{\mathbf{curl}_{\mathrm{h}}}=0,\\ &\int_{{r{\mathbb{S}^{2}}}}\operatorname{\mathbf{grad}_{\mathrm{h}}}f\cdot\textbf{{f}}\,\mathrm{d}s=-\int_{{r{\mathbb{S}^{2}}}}f\operatorname{div_{\mathrm{h}}}\textbf{{f}}\,\mathrm{d}s,\,\,\,\int_{{r{\mathbb{S}^{2}}}}\operatorname{\mathbf{curl}_{\mathrm{h}}}f\cdot\textbf{{f}}\,\mathrm{d}s=\int_{{r{\mathbb{S}^{2}}}}f\operatorname{curl_{\mathrm{h}}}\textbf{{f}}\,\mathrm{d}s.\end{split}$ (50)
@@
and $\langle\operatorname{\mathbf{grad}_{\mathrm{h}}}\varphi,\operatorname{\mathbf{curl}_{\mathrm{h}}}\psi\rangle_{L^{2}}=0$ (see [28, Eq. (5.6.24)]).
@@
$H^{s}({r{\mathbb{S}^{2}}}):=\mathrm{dom}((I-\Delta_{\rm h})^{s/2})\qquad\mbox{with}\quad\|f\|_{H^{s}}:=\|(I-\Delta_{\rm h})^{s/2}f\|_{L^{2}}$
@@
$L^{2}_{\diamond}({r{\mathbb{S}^{2}}}):=\{f\in L^{2}({r{\mathbb{S}^{2}}}):\langle f,1\rangle=0\},\quad H^{s}_{\diamond}({r{\mathbb{S}^{2}}}):=\{f\in H^{s}:\langle f,1\rangle=0\}$ (54)
@@
The following embeddings hold true (see [36, Propositions 3.2-3.3], [17, Theorem 3.5]):
@@
$L^{2}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m=-\infty}^{\infty}L^{2}({r{\mathbb{S}^{2}}};m),\quad L^{2}_{\diamond}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m=-\infty}^{\infty}L^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)$ (57)
@@
$L^{2}({r{\mathbb{S}^{2}}};m)=L^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)=\overline{\mathrm{span}\{Y_{lm}:l\geq|m|\}}^{L^{2}},\qquad m\in\mathbb{Z}\setminus\{0\},$
$L^{2}({r{\mathbb{S}^{2}}};0)=\overline{\mathrm{span}\{Y_{l0}:l\in\mathbb{Z}\}}^{L^{2}},\quad L^{2}_{\diamond}({r{\mathbb{S}^{2}}};0)=\overline{\mathrm{span}\{Y_{l0}:l\in\mathbb{Z}\setminus\{0\}\}}^{L^{2}},$
@@
$\Delta_{\mathrm{h}}f=\sum_{m=-\infty}^{\infty}(\Delta_{m}\hat{f}_{m})\otimes e_{m}\quad\mbox{with}\quad\Delta_{m}:=\frac{1}{r^{2}\sin\theta}\frac{d}{d\theta}\left(\sin\theta\frac{d}{d\theta}\right)-\frac{m^{2}}{r^{2}\sin^{2}\theta}\,$ (58)
@@
$H^{s}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m\in\mathbb{Z}}H^{s}({r{\mathbb{S}^{2}}};m)\quad\mbox{and}\quad H^{s}_{\diamond}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m\in\mathbb{Z}}H_{\diamond}^{s}({r{\mathbb{S}^{2}}};m)$ (60)
@@
$\|f\|_{H_{\diamond}^{s}}:=\|(-\Delta_{\mathrm{h}})^{s/2}f\|_{L^{2}},\qquad s\geq 0$
@@
$\begin{aligned} &\|f\|_{H_{\diamond}^{2}}=\|\Delta_{\mathrm{h}}f\|_{L^{2}}\quad\mbox{and}\quad\\ &\|f\|_{H_{\diamond}^{1}}^{2}=\langle-\Delta_{\mathrm{h}}f,f\rangle=\langle-\operatorname{div_{\mathrm{h}}}\operatorname{\mathbf{grad}_{\mathrm{h}}}f,f\rangle=\|\operatorname{\mathbf{grad}_{\mathrm{h}}}f\|_{L^{2}}^{2}.\end{aligned}$ (61)
@@
$\Delta_{\mathrm{h}}:H^{t}_{\diamond}({r{\mathbb{S}^{2}}})\to H^{t-2}_{\diamond}({r{\mathbb{S}^{2}}})\quad\text{is an isometric isomorphism for all $t\in\mathbb{R}$.}$ (62)

### v3

@@ 5.1 Full measurement
With the TCC, we further obtain unique identifiability of $\Omega$ given $\gamma$ and vice versa. We denote by $\mathrm{Arg}(z)\in[0,2\pi)$ the principal value of the argument of a complex number $z$.
@@
    $\Delta_{m}\Psi\neq 0\text{ a.e.~on }I,\qquad\mathrm{Arg}(\Psi)-\mathrm{Arg}(\Delta_{m}\Psi)\notin\{-\pi,0,\pi\}\text{ a.e.~on }I.$ (47)
@@
From this, along with the characterization of the solution set given the TCC in [HNS95, Proposition 2.1], we deduce $\Omega^{*}-\Omega^{\dagger}\in\mathcal{N}(F^{\prime}(\Omega^{\dagger}))=\{0\}$ for any solution $\Omega^{*}$, $\Omega^{\dagger}$ of the inverse problem. We thus claim the unique identifiability of $\Omega$ given full measurement.
@@ 5.2 Partial measurement including Cauchy data
In the last part of the analysis, we consider the restricted measurement scenario where observational latitude range is limited to $I^{\prime}=(\epsilon,\pi-\epsilon)$ as described in Section 2.3. For simplification, we additionally assume that Cauchy data at the boundary of $I^{\prime}$ can be measured exactly. We follow the analysis developed for the full observation, but specifically leveraging the well-posedness and regularity results for the restricted domain from Proposition 9 in Section 3.5.
@@
$\Psi|_{\partial I^{\prime}}=\Psi^{\dagger}|_{\partial I^{\prime}}\qquad\text{and}\qquad\Psi^{\prime}|_{\partial I^{\prime}}=(\Psi^{\dagger})^{\prime}|_{\partial I^{\prime}},$ (49)
@@
The boundary data assumption in (49), requiring knowledge of the exact state at $\partial I^{\prime}$, is important for the proof of Theorem 17. Relaxing this condition presents an interesting question for future research. With this theoretical discussion concluded, we now turn to the numerical experiments.
@@ 6 Numerical experiments
This section is dedicated to numerical experiments on synthetic data, using analytical ground truths. We implemented the accelerated Nesterov-Landweber method to simultaneously reconstruct the scalar viscosity and latitudinal differential rotation parameters $(\gamma,\Omega)$. A convergence analysis of this method has been carried out in [HubmerRamlau17, Neubauer17] under the tangential cone condition, which we have been verified in Section 5. The Nesterov-Landweber iteration proceeds as follows: initiate $p_{-1}=p_{0}:=(\gamma_{\mathrm{int}},\Omega_{\mathrm{int}})$, then run
$z_{k}$ $=p_{k}+\frac{k-1}{k+\alpha-1}\left(p_{k}-p_{k-1}\right)\hskip 18.49988pt$ $\alpha=3$ (50)
@@
Here, the forward map $F$ corresponds to the wave equation at single frequency $\omega$ and longitudinal wavenumber $m$. We allow for different observation strategies: full data, leaked data, and real part measurements. The Hilbert space adjoints $F^{\prime}(z_{k})^{*}$ are derived in Section 4.2. The inversion takes place in the function spaces: $\gamma\in\mathbb{R}$, $\Omega\in H^{2}({r{\mathbb{S}^{2}}},m=0)$, ensuring convergence of the iterative reconstruction.
The numerical wave solver uses a fourth-order finite difference scheme [VersyptBraatz14] on a uniform grid of 100 points spanning the latitude range $(0,\pi)$. To generate ground truths while avoiding inverse crimes, we choose analytic $\Psi_{m}$$\gamma$, and $\Omega$ with Neumann (Fig. 1) or Dirichlet (Fig. 2) boundary, then explicitly obtain the corresponding source $f$. The discrete measurement $\underline{y}^{\delta}$ is corrupted by independent additive complex Gaussian noise with variance $\sigma^{2}$ on 100 equidistant measurement points and $\sigma$ chosen to have relative noise levels $\|\underline{y}^{\delta}-\underline{y}\|_{2}/\|\underline{y}\|_{2}\in\{0.01,0.05,0.1,0.2\}$.
We initialize the algorithm with $\gamma_{\mathrm{int}}=-3\gamma_{\mathrm{true}}$, $\Omega_{\mathrm{int}}=0$ corresponding to no prior knowledge of ground truths, and iterate with a step size $\mu_{k}$ determined by backtracking line search. Following a discrepancy principal, the stopping criterion is the first instance when the residual falls below the threshold $\tau\delta$ with $\tau=1.03$; this in our practice yields better results than the theoretical $\tau>2$ in [HubmerRamlau17]. The algorithm typically completes 200 iterations within one second on an i7-1255U CPU (4.70 GHz) [^1] Code is available on Github at https://github.com/TramNguyenAca/Inertial_Waves..
@@
Figure 1, row 1 shows the simultaneous recovery of $(\gamma,\Omega)$ from full measurement of the state $\Psi$ at $(\omega,m)=(3,3)$ and with $1\%$ data noise. Despite of an initial guess far from the ground truth, the reconstruction closely approximates the exact parameters. The approximated wave field also matches the true state in both real and imaginary parts.
@@
Figure 1, rows 2–3 present the practical case where no data are available near the poles. In our numerical experiments, we did not impose additional Cauchy boundary data in Theorem 17, and simply computed the adjoint as described in Remark 13. With 20% of the data leaked (row 2), the reconstruction quality remains nearly as high as with full data. For 50% data loss (row 3), the recovery of $\Omega$ moderately deteriorates. Nonetheless, the algorithm exhibits robust performance, consistent with our convergence results in Section 5.2.
@@
Figure 1, row 4 presents the case where only the real part of $\Psi$ can be measured. Compared to row 3 (which also has 50% data leakage, but full complex data), the absence of the imaginary component further deteriorates the reconstruction.
@@
Finally, Figure 2 reports the results for $(\omega,m)=(1,2)$ with 20 % data leakage and different noise levels. At 1% noise (row 1), the recovered parameters remain very close to the truth, as evidenced by small relative errors ($\||p-p_{\text{true}}\|_{L^{2}}/\|p_{\text{true}}\|_{L^{2}}$). At 5% noise (row 2), reconstruction quality is slightly reduced, while at 20% noise (row 4), the outcome degrades significantly.
@@ 7 Conclusions and Outlook
In this paper, we investigated the modeling, as well as the associated forward and inverse problems for inertial waves on the surface of the Sun. Formulating the dynamics in terms of a stream function leads to a fourth-order elliptic differential equation on the sphere and its separated version for each longitudinal-wavenumber. We rigorously proved well-posedness of these equations, including explicit smallness conditions for uniqueness, as well as conditions for non-uniqueness via analytic Fredholm theory, both in agreement with empirical findings in helioseismology. Our studies represent the starting point for further theoretical and numerical studies of resonances and inertial modes, as well as the associated inverse problems.
@@
Future extensions of this work include incorporating realistic stochastic sources of solar excitation [PhilidetGizon23] within a passive imaging framework. Additionally, we will investigate the nonlinear equation for finite amplitude waves, which is required to interpret the time evolution of linearly unstable inertial modes [Bekki2024]. Another promising avenue involves leveraging data-driven model discovery techniques [AarsetHollerNguyen23, nguyen25] to refine existing theoretical inertial wave models by learning hidden laws.
@@ Appendix A Differential geometry and function spaces on spheres
independent of the choices of $\tilde{f}$ and $\tilde{\textbf{{f}}}$ (see [nedelec:01, p. 72–73]). Here $\operatorname{div_{\mathrm{h}}}\textbf{{f}}$ and $\operatorname{curl_{\mathrm{h}}}\textbf{{f}}$ are scalar functions, while $\operatorname{\mathbf{grad}_{\mathrm{h}}}f$ and $\operatorname{\mathbf{curl}_{\mathrm{h}}}v$ are tangential vector fields on $r{\mathbb{S}^{2}}$}. In spherical coordinates, writing $\textbf{{f}}=f_{\theta}\mathbf{e}_{\theta}+f_{\phi}\mathbf{e}_{\phi}$ with $f_{\theta}:=\textbf{{f}}\cdot\mathbf{e}_{\theta}$ and $f_{\phi}:=\textbf{{f}}\cdot\mathbf{e}_{\phi}$, we have
@@
These differential operators satisfy the identities ([nedelec:01, Thm. 2.5.19]):
$\begin{split}&\operatorname{curl_{\mathrm{h}}}\operatorname{\mathbf{grad}_{\mathrm{h}}}=0,\hskip 18.49988pt\operatorname{div_{\mathrm{h}}}\operatorname{\mathbf{curl}_{\mathrm{h}}}=0,\\ &\int_{{r{\mathbb{S}^{2}}}}\operatorname{\mathbf{grad}_{\mathrm{h}}}f\cdot\textbf{{f}}\,\mathrm{d}s=-\int_{{r{\mathbb{S}^{2}}}}f\operatorname{div_{\mathrm{h}}}\textbf{{f}}\,\mathrm{d}s,\,\,\,\int_{{r{\mathbb{S}^{2}}}}\operatorname{\mathbf{curl}_{\mathrm{h}}}f\cdot\textbf{{f}}\,\mathrm{d}s=\int_{{r{\mathbb{S}^{2}}}}f\operatorname{curl_{\mathrm{h}}}\textbf{{f}}\,\mathrm{d}s.\end{split}$ (53)
@@
and $\langle\operatorname{\mathbf{grad}_{\mathrm{h}}}\varphi,\operatorname{\mathbf{curl}_{\mathrm{h}}}\psi\rangle_{L^{2}}=0$ (see [nedelec:01, Eq. (5.6.24)]).
@@
$H^{s}({r{\mathbb{S}^{2}}}):=\mathrm{dom}((I-\Delta_{\rm h})^{s/2})\hskip 18.49988pt\mbox{with}\qquad\|f\|_{H^{s}}:=\|(I-\Delta_{\rm h})^{s/2}f\|_{L^{2}}$
@@
$L^{2}_{\diamond}({r{\mathbb{S}^{2}}}):=\{f\in L^{2}({r{\mathbb{S}^{2}}}):\langle f,1\rangle=0\},\qquad H^{s}_{\diamond}({r{\mathbb{S}^{2}}}):=\{f\in H^{s}:\langle f,1\rangle=0\}$ (57)
@@
The following embeddings hold true (see [Taylor, Propositions 3.2-3.3], [Hebey, Theorem 3.5]):
@@
$L^{2}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m=-\infty}^{\infty}L^{2}({r{\mathbb{S}^{2}}};m),\qquad L^{2}_{\diamond}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m=-\infty}^{\infty}L^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)$ (60)
@@
$L^{2}({r{\mathbb{S}^{2}}};m)=L^{2}_{\diamond}({r{\mathbb{S}^{2}}};m)=\overline{\mathrm{span}\{Y_{lm}:l\geq|m|\}}^{L^{2}},\hskip 18.49988ptm\in\mathbb{Z}\setminus\{0\},$
$L^{2}({r{\mathbb{S}^{2}}};0)=\overline{\mathrm{span}\{Y_{l0}:l\in\mathbb{Z}\}}^{L^{2}},\qquad L^{2}_{\diamond}({r{\mathbb{S}^{2}}};0)=\overline{\mathrm{span}\{Y_{l0}:l\in\mathbb{Z}\setminus\{0\}\}}^{L^{2}},$
@@
$\Delta_{\mathrm{h}}f=\sum_{m=-\infty}^{\infty}(\Delta_{m}\hat{f}_{m})\otimes e_{m}\qquad\mbox{with}\qquad\Delta_{m}:=\frac{1}{r^{2}\sin\theta}\frac{d}{d\theta}\left(\sin\theta\frac{d}{d\theta}\right)-\frac{m^{2}}{r^{2}\sin^{2}\theta}\,$ (61)
@@
$H^{s}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m\in\mathbb{Z}}H^{s}({r{\mathbb{S}^{2}}};m)\qquad\mbox{and}\qquad H^{s}_{\diamond}({r{\mathbb{S}^{2}}})=\bigoplus\nolimits_{m\in\mathbb{Z}}H_{\diamond}^{s}({r{\mathbb{S}^{2}}};m)$ (63)
@@
$\|f\|_{H_{\diamond}^{s}}:=\|(-\Delta_{\mathrm{h}})^{s/2}f\|_{L^{2}},\hskip 18.49988pts\geq 0$
@@
$\begin{aligned} &\|f\|_{H_{\diamond}^{2}}=\|\Delta_{\mathrm{h}}f\|_{L^{2}}\qquad\mbox{and}\qquad\\ &\|f\|_{H_{\diamond}^{1}}^{2}=\langle-\Delta_{\mathrm{h}}f,f\rangle=\langle-\operatorname{div_{\mathrm{h}}}\operatorname{\mathbf{grad}_{\mathrm{h}}}f,f\rangle=\|\operatorname{\mathbf{grad}_{\mathrm{h}}}f\|_{L^{2}}^{2}.\end{aligned}$ (64)
@@
$\Delta_{\mathrm{h}}:H^{t}_{\diamond}({r{\mathbb{S}^{2}}})\to H^{t-2}_{\diamond}({r{\mathbb{S}^{2}}})\qquad\text{is an isometric isomorphism for all $t\in\mathbb{R}$.}$ (65)